OpenStudy (anonymous):

A simple pendulum has a period of 1.870 seconds. What is the length of the pendulum?

OpenStudy (anonymous):

T = 1.870 g = 9.8 \[T = 2\pi \sqrt{\frac{ L }{ g }}\] rearrange for L \[L = \frac{ T^2g }{ 4\pi^2 }\] L = 86.805cm

OpenStudy (anonymous):

Thanks so so much!